The Friendly Sausage Factory (FSF) can produce hot dogs at a rate of 500 per day. FSF supplies hot dogs to local restaurants at a steady rate of 250 per day (open 300 days). The cost to prepare the equipment for producing hot dogs is $30. Annual holding costs are $2 per hot dog. The factory operates 300 days a year.

  1. Evaluate the number of computers that the store manager should order in each replacement lot.
  2. Maximum Inventory
  3. Average Inventory
  4. Length of the order cycle
  5. Annual inventory-related cost (Yearly total cost)
Answer #1

Production Rate p = 500 per day

Daily Demand d = 250

Annual Demand D = 250*300 = 75000

Set up Cost S = 30

Holding Cost H = 2

a.

Economic Production Quantity EPQ = (2DS/(H*(1-(d/p))))^(1/2)

EPQ = (2*75000*30/(2*(1-(250/500))))^(1/2)

EPQ = 2121.32 units

b.

Imax = EPQ*(1-(d/p))

Imax = 2121.32*(1-(250/500))

Imax = 1060.66 units

c.

Average Inventory = Imax/2 = 1060.66/2 = 530.33

d.

Order Cycle Length = EPQ/d = 2121.32/250 = 8.49 days

Production run time = EPQ/p = 2121.32/500 = 4.24 days

e.

Annual Inventory Cost = (D/EPQ)*S + (Imax/2)*H

Annual Inventory Cost = (75000/2121.32)*30 + (2121.32/2)*2 = 3181.98